Circular No. 7674

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dj IN NGC 180 M. Papenkova and W. D. Li, University of California, report the dis-

covery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on July 29.5 (mag ∼ 17.9) and 30.5 UT (mag ∼ 17.7). The new object is located at α = 0^{h}37^{m}52^{s}.22, δ = +8^{o}38′33′′.0 (equinox 2000.0), which is 80′′.8 west and 27′′.1 north of the nucleus of NGC 180. A KAIT image taken on Jan. 5.2 showed nothing at this position (limiting mag ∼ 19.0).

NOVA IN M31 Li also report the discovery of an apparent nova in M31 (= NGC 224)

on unfiltered KAIT images taken on July 28.5 (mag ∼ 18.0) and 30.5 UT (mag ∼ 17.5). The nova is located at α = 0^{h}42^{m}30^{s}.72, δ = +41^{o}14′36′′.8 (equinox 2000.0), which is 153′′ west and 92′′ south of the nucleus of M31. A KAIT image of the same field taken on July 26.5 showed nothing at this position (limiting mag ∼ 19.0).

V1494 AQUILAE R. J. Rudy, S. Mazuk, C. Venturini, and D. K. Lynch, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and P. Mortfield, Stanford University, report on 0.8–2.5-μm spectrophotometry of V1494 Aql, using the Lick Observatory 3-m telescope (+ Aerospace Near Infrared Imaging Spectrograph) on July 10.6 UT, ∼ 580 days after peak brightness: “The nova is in a strengthening coronal phase (see IAUC 7490) with lines of [Al IX], [Ca VIII], [S VIII], [S IX], [Si VI], [Si VII], [S IX], [Si X], and possibly [S XI] present. The latter, at 1.9196 μm, requires photon energies > 447 eV to produce. Many of the lines have double-peaked profiles and have typical line widths (full width at half maximum) of 2900 km/s. The continuum decreased monotonically with increasing wavelength and showed no evidence of thermal emission from dust.”

COMETS C/2000 S3, 150P, C/2000 Y2, C/2001 G1, C/2001 HT_{50} The IAU Committee on Small Bodies Nomenclature has agreed upon

the names for the following five comets: C/2000 S3 (LONEOS); 150P/2000 WT_{168} (LONEOS); C/2000 Y2 (Skiff); C/2001 G1 (LONEOS); C/2001 HT_{50} (LINEAR-NEAT).

2001 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green